If X has a normal distribution with mean µ = 12 and variance σ2 = 9, compute P (12 < X < 15).

A

0.3413

B

0.9651

C

0.8413

D

0.5000
Answer #1

Solution :

Given that ,

mean = = 12

standard deviation = = 3

P(12 < x < 15) = P[()12 - 12/ 3) < (x - ) /  < (15 - 12) / 3) ]

= P(0 < z < 1)

= P(z < 1) - P(z < 0)

= 0.8413 - 0.5

= 0.3413

P(12 < x < 15) = 0.3413

option A is correct